[4 years ago Shak was 5 times as old as his younger brother John. If John will be half as old as Shak in 2 years, how old is John?]
Answer by TimothyLamb(4379)

x = john age now
y = shak age now
---
y - 4 = 5(x - 4)
y + 2 = 2(x + 2)
---
y - 4 = 5x - 20
y + 2 = 2x + 4
---
5x - y = 16
2x - y = -2
---
put the system of linear equations into standard form
---
5x - y = 16
2x - y = -2
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = john age now = 6
y = shak age now = 14
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php